OpenStudy (aaronandyson):

The age of father was twice the square of the age of his son.Eight years,hence,The age of the father will be 4 years more than 3 times the age of his son.Find their present ages.

OpenStudy (kropot72):

Should this question read as follows: "The age of father is twice the square of the age of his son. Eight years hence, the age of father will be 4 years more than 3 times the age of his son. Find their present ages."

OpenStudy (aaronandyson):

@Sepeario

OpenStudy (sepeario):

Yes @kropot72 the question doesn't actually make sense otherwise.

OpenStudy (sepeario):

If the age of the father is x and the age of the son is y, then we get the simultaneous equations: \[x=2y^2\] and also \[x+8=4+3(y+8)\]

OpenStudy (sepeario):

I think..

OpenStudy (sepeario):

If you rearrange the second equation you get \[x=3y+20\] and then you substitute the first equation into the second, leaving \[2y^2=3y+20\] or \[2y^2-3y-20=0\]

OpenStudy (sepeario):

You can then factorise this polynomial equation.

OpenStudy (sepeario):

Giving (2y+5)(y-4), leaving the solutions for y at either 4 or -5/2

OpenStudy (aaronandyson):

4 and 32

OpenStudy (sepeario):

There you go!

OpenStudy (sepeario):

Yes exactly.